ShadowKid3:

5(−2x−1)−5x+2=27 solve for x

luhivqqcherry:

5(-2x-1)-5x+2=27. x=27

ShadowKid3:

nope

luhivqqcherry:

x=-2

luhivqqcherry:

i didn't read the question right the first time.

OLIVER69:

Problem: 5(−2x−1)−5x+2=27 Step 1: Simplify both sides of the equation. 5(−2x−1)−5x+2=27 (5)(−2x)+(5)(−1)+−5x+2=27(Distribute) −10x+−5+−5x+2=27 (−10x+−5x)+(−5+2)=27(Combine Like Terms) −15x+−3=27 −15x−3=27 Step 2: Add 3 to both sides. −15x−3+3=27+3 −15x=30 Step 3: Divide both sides by -15. −15x/−15=30/−15 x=−2 Answer: x=−2

OLIVER69:

@luhivqqcherry wrote:
x=-2
Stop giving direct answers

luhivqqcherry:

@oliver69 wrote:
@luhivqqcherry wrote:
x=-2
Stop giving direct answers
why

OLIVER69:

@luhivqqcherry wrote:
@oliver69 wrote:
@luhivqqcherry wrote:
x=-2
Stop giving direct answers
why
Because it doesn't help anyone learn... plus you will get reported that way.

luhivqqcherry:

@oliver69 wrote:
@luhivqqcherry wrote:
@oliver69 wrote:
@luhivqqcherry wrote:
x=-2
Stop giving direct answers
why
Because it doesn't help anyone learn... plus you will get reported that way.
mhm

hannahj:

@luhivqqcherry wrote:
@oliver69 wrote:
@luhivqqcherry wrote:
@oliver69 wrote:
@luhivqqcherry wrote:
x=-2
Stop giving direct answers
why
Because it doesn’t help anyone learn... plus you will get reported that way.
mhm
lol 😆

HotPockets14:

only if you report her see will be. I dont see what the problem is

HotPockets14:

she

luhivqqcherry:

exactly its only an answer .

luhivqqcherry:

hes pressed for no reason.

HotPockets14:

yep

Luigi0210:

Not necessarily pressed, I think they're just trying to follow the rules lol

HotPockets14:

@luigi0210 wrote:
Not necessarily pressed, I think they're just trying to follow the rules lol
Okay but a question, why do they have zero tolerance for direct answers. Im just wondering

Luigi0210:

It's a homework HELP site, not a homework answer bank site. It was created with the intention of helping students and others with the learning process, aka guiding them to answer rather than telling them how to think.

Luigi0210:

Or you can just ask our site overlord @ultrilliam pretty sure he's the creator but I'm not 100% sure

hannahj:

@luigi0210 wrote:
It’s a homework HELP site, not a homework answer bank site. It was created with the intention of helping students and others with the learning process, aka guiding them to answer rather than telling them how to think.
yes yes 🙌🏼

HotPockets14:

@hannahj wrote:
@luigi0210 wrote:
It’s a homework HELP site, not a homework answer bank site. It was created with the intention of helping students and others with the learning process, aka guiding them to answer rather than telling them how to think.
yes yes 🙌🏼
okay i was just wondering

HotPockets14:

@hannahj wrote:
@luigi0210 wrote:
It’s a homework HELP site, not a homework answer bank site. It was created with the intention of helping students and others with the learning process, aka guiding them to answer rather than telling them how to think.
yes yes 🙌🏼
okay i was just wondering

Veequeen:

x =−2